OpenStudy (anonymous):

solove the equation and check the solution x-21=-31

OpenStudy (nincompoop):

let me see you attempt it first

OpenStudy (anonymous):

i need help with it thats why i am posting it. i got it wrong on my test.

OpenStudy (nincompoop):

how about you study these examples from a very simple 4 - 5 = -1 4 -x = -1 4 - 4 - x = -1 -4 -x = -5 x = 5 see how x = 5 makes 4 - x = -1 how about another one 3 + y = 7 3 - 3 + y = 7 - 3 y = 4 3 + 4 = 7

OpenStudy (nincompoop):

the first process is to isolate or leave the variable that you are solving on one side by using the opposite signs of the terms it was with. whatever subtraction you will perform on the one side, must be done on the other side of the equation, the same goes for addition, multiplication and division

OpenStudy (nincompoop):

thank you for your time and I hope you learned something new tonight. and there is no need to thank me.